Solve the system by the elimination method.
5x-2y=3
10x-4y=5

Answer by mananth(16946)

5 x -2 y = 3 .............1
Total value
10 x -4 y = 5 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 2
-20 x 8 y = -12
20 x -8 y = 10
Add the two equations
0 x = -2

NO SOLUTION
The lines are parallel